is the lines through points P(4,-7)and Q(3,-4) parallel to the line through points R(-5,3) And S(-2,2)?
Answer by mananth(16946)

slope m = (y2-y1)/(x2-x1)
P(4,-7)and Q(3,-4)

( -4 - -7 )/( 3 - 4 )
( 3 / -1 )
m= -3



R(-5,3) And S(-2,2)
x1 y1 x2 y2
-5 3 -2 2

slope m = (y2-y1)/(x2-x1)
( 2 - 3 )/( -2 - -5 )
( -1 / 3 )
m= - 1/3
The slopes are not equal so they are not parallel